A corporation issues 1,200,000 shares of stock at its beginning to shareholders. How many shares must
a shareholder own to have a majority of the shares?

Found 2 solutions by jim_thompson5910, stanbon:
Answer by jim_thompson5910(35256)

A shareholder must have 50% + 1 share to have the majority

50% of 1,200,000 = 0.50*1200000 = 600,000

add one more share to get over the 50% threshold: 600,000 + 1 = 600,001

So a shareholder must own 600,001 shares

Answer by stanbon(75887)

A corporation issues 1,200,000 shares of stock at its beginning to shareholders. How many shares must
a shareholder own to have a majority of the shares?
-----
Find the half-way point:
1200000/2 = 600,000
----
A majority share is at least 600,001
========================================
Cheers,
Stan H.
================